A freeway ramp is 10 ft high after 200 ft. How high is the highest point of the ramp if the ramp is 1500 ft in length?

Answer by Theo(13342)

the slope is the change in the vertical height divided by the corresponding change in the horizontal distance.
10/200 = 1/20.
1/20 * 1500 = 75 feet high.
the ramp will have a vertical height of 75 feet after a horizontal distance of 1500 feet, assuming it continues ramping up at the same rate as when it had a vertical height of 10 feet after a horizontal distance of 200 feet.